# Sequential Regression Learning with Randomized Algorithms

### v1

@@ 4. Randomized SINDy
$\ell(y,\mathbf{x}^{\top}w)=-\bigr{[}y\ln\left(h\left(\mathbf{x}\right)\right)+\left(1-y\right)\ln\left(1-h\left(\mathbf{x}\right)\right)\bigr{]},$
we note that there is no closed form for $\mathbb{E}_{W\sim\mathcal{N}(\mu,\Sigma)}\ell(y,\mathbf{x}^{\top}w)$. Rather, we consider its approximation $\ell(y,\mathbf{x}^{\top}\mu)$. Consequently, the objective function is as follows:
$\underbrace{\ell(y,\mathbf{x}^{\top}\mu)+\lambda\left(\text{tr}\Sigma+\mu^{\top}\mu\right)}_{f_{1}}+\underbrace{1\!\!1_{p.s.d}(\Sigma)}_{f_{2}}$
@@
$\nabla_{\mu}f_{1}(\mu,\Sigma)=\mathbf{x}\left(\sigma\left(\mathbf{x}^{\top}\mu\right)-y\right)+2\lambda\mu,$
@@
$\mu_{t+1}=\mu_{t}-\eta_{t}\nabla_{\mu_{t}}f_{1}(\mu_{t},\Sigma_{t})=\mu_{t}-\eta_{t}\left(\mathbf{x}_{t}\left(\frac{1}{1+e^{-\mathbf{x}^{\top}_{t}\mu_{t}}}-y_{t}\right)+2\lambda\mu_{t}\right),$
@@
5. Randomized Algorithm: Finite-dimensional Euclidean Space
Considering the case of a finite predictor space $\mathcal{H}$, the learner has access to $N$ reference predictors, commonly referred to as experts in the context of online learning, at every time step. Each expert provides a prediction $h_{i}(\mathbf{x})$ for a given input vector $\mathbf{x}\in\mathbb{R}^{k}$. We define a discrete measure $\delta$ over =$\mathcal{H}=\{h_{1},h_{2},\cdots,h_{N}\}$, such that each expert $h_{i}$ is assigned a probability $g_{i}$. The probability distribution over $\mathcal{H}$ is represented by the vector:
$\Delta^{N}=\left\{g=(g_{1},\cdots,g_{N}):0\leq g_{i}\leq 1,\text{for }i=1,\cdots,N\text{ and}\sum_{i}^{N}g_{i}=1\right\}.$
Here, $g_{i}=P\left(h_{i}\left(\mathbf{x}\right)\right)$ indicates the weight assigned to expert $i$’s prediction, and the set of all such vectors $g$ forms the standard $N$-simplex, denoted by $\Delta^{N}$.
At each time step $t$, after observing the pair $(\mathbf{x}_{t},y_{t})$, $\ell(y_{t},h_{i}\left(\mathbf{x_{t}}\right))$ denotes the loss incurred when expert $i$ predicts $h_{i}$ based on $\mathbf{x_{t}}$. The expected loss under the current distribution $g_{t}$ is:
$c_{(\mathbf{x}_{t},y_{t})}(g_{t})=\sum_{i=1}^{N}\ell(y_{t},h_{i}(\mathbf{x}_{t}))g_{t}(h_{i}).$
We initialize $g_{1}:=1$. To update the probability distribution for $t\geq 1$, we first compute an intermediate vector using a gradient step:
$z_{t}(h_{i})$ $=$ $g_{t}(h_{i})-\eta_{t}\nabla c_{(\mathbf{x}_{t},y_{t})}(g_{t})$
$=$ $g_{t}(h_{i})-\eta_{t}\ell(y_{t},h_{i}(\mathbf{x}_{t}))$
where $\eta_{t}$ is the learning rate. Since vector $z_{t}$ might not be a valid probability distribution as it may contain negative values or not sum to 1, we project it onto the simplex $\Delta_{N}$ using the proximal operator with an indicator function $\Psi$:
$g_{t+1}$ $=$ $\text{prox}_{\eta_{t}\Psi}(z_{t})$
$=$ $\text{proj}_{\Delta^{N}}(z_{t}).$
This projection ensures that the updated distribution $g_{t+1}$ remains a valid probability distribution over the set of experts $h_{i}$.
@@ 6. Experiments
In this section, we present some simulation results of models discussed in section 4 and section 5. All experiments were executed using the R language.
@@
To evaluate the impact of the initial value, we vary $\mu_{0}$ between 0 and 1.8. As shown in Figure 2, the initial values influence the convergence rate, as stated in Lemma 3.1. For a sequence observed at the 50th time step, the initial value of $\mu_{0}=0$ requires more iterations to achieve convergence.
@@
The first 250 time steps were designated as a warm-up period for model estimation. Within this warm-up period, we performed a rolling window cross-validation with 40 train-validation splits. The first training set consisted of 50 observations. For each subsequent split, the training set size was incremented by five observations (e.g.. 50, 55, …, up to 245 observations). Each corresponding validation set contained the five subsequent observations immediately following its training set. For each of these 40 train-validation splits, we fitted a regularized regression model. The regularization parameter ($\lambda$) was varied across 30 different values, ranging from 0 to 1. To determine the optimal $\lambda$, the minimum Root Mean Squared Error (RMSE) was calculated for each $\lambda$ value on each of the 40 validation sets. Subsequently, $\lambda=0.3448$ was selected as the optimal $\lambda$ as it yielded the lowest overall RMSE. Still within this warm-up period, we proceeded to standardize the model matrix, which has dimensions of $250\times 4$. A preliminary Tikhonov regularization was then performed to obtain an initial guess for the mean coefficient vector ($\mu_{0}$) and its covariance matrix ($\Sigma_{0})$. Subsequently, the algorithm was applied with the parameter update steps defined by Equations 4.4 and 4.5. The model parameters estimated at $t=250$ were used as the initial state for the subsequent time step. From $t=251$ onwards, each new incoming data dynamically updated both the standardization of the model matrix and the estimated parameters. With a learning rate of $\eta=1\times 10^{-3}$, Table 3 presents the estimated parameters at the initial and the final time step.
@@
$t$ $SST$ $SSE$ $n$ $p$ $R^{2}$ $\hat{\sigma}$ $RMSE$
@@
We simulated a binary sequence $Y\sim Ber(p)$, where $p$ was defined by the logistic model $P(Y=1\mid X)=\dfrac{1}{1+\exp\{-(1+2X_{1}+3X_{2})\}}.$ As in the previous experiment, a sequence of 500 time steps was generated. Initial logistic model parameters were obtained by fitting a ridge logistic regression with $\lambda=0$. The algorithm from Example 2 of section 4 was then applied with a fixed learning rate of $\eta=0.1$ Table 5 shows the estimated coefficients at both the initial and final time steps.
@@
To evaluate the ensemble method described in section 5, we simulated a sequence of a size of 100 where each outcome was generated based on the following model:
$Y=1+1.1X_{1}+1.2X_{2}+1.3X_{3}+1.4X_{4}+\epsilon,$
where $\epsilon\sim\mathcal{N}(0,0.5)$, and $X_{i}\sim\mathcal{U}(0,1)$. The hypothesis space consisted ten distinct expert models, as detailed in Table 7. It is evident that the model from expert 10 most closely approximates to the data generating process. Consequently, we expect that the algorithm will converge by assigning the majority of the weight to this expert.
Table 7. Hypothesis space of experts
Expert Model Expert Model
$E_{1}$ $Y=X_{1}$ $E_{6}$ $Y=1+1.2X_{2}$
$E_{2}$ $Y=X_{2}$ $E_{7}$ $Y=1+1.3X_{3}$
$E_{3}$ $Y=X_{3}$ $E_{8}$ $Y=1+1.4X_{4}$
$E_{4}$ $Y=X_{4}$ $E_{9}$ $Y=1+1.1X_{1}+1.2X_{2}+1.3X_{3}$
$E_{5}$ $Y=1+1.1X_{1}$ $E_{10}$ $Y=1+1.1X_{1}+1.2X_{2}+1.3X_{3}+1.4X_{4}$
Using a learning rate of $\eta=0.1$, we evaluated two prediction methods: one based on weighted experts and another on the best expert. The outcomes, depicted in Figures 4 and 5, reveal that the weight assignments for both methods converged to expert 10, with its weight increasing to 1 after 50 time steps. The prediction method based on the best expert in this case consistently yielded a lower cumulative loss.
Figure 4. Evolution of weight assignments to each expert
Figure 5. Cumulative loss of each expert along $t$
Experiment 5
We applied the models discussed in section 4 to forecast the monthly U.S. unemployment rate. As described in the work of [Elec2], the independent variables included are Initial Claims (IC) for unemployment insurance, the inflation rate (measured as the seasonal growth rate of the Consumer Price Index), and industrial growth (the seasonal growth rate of the Industrial Production Index). All data were sourced from the FRED database (https://fred.stlouisfed.org/). The analysis period spans from January 1967 to May 2025. Table 8 gives a brief description of the variables used for forecasting. While most variables are originally monthly, the weekly Initial Claims data was averaged to a monthly frequency.
@@
We used the initial 30 years of data, spanning from January 1967 to December 1999, as a warm-up period for our online learning model. Based on a rolling expanding window approach, we began with 20 years of data for training to make a one-year-ahead prediction. This process generated 12 distinct train-validation set splits for model evaluation. For each of these 12 training sets, we performed regression with Tikhonov regularization, exploring 30 different regularization parameter ($\lambda$) values ranging from 0 to 1. The optimal regularization term was identified as being the value of 0, as this produced the minimum overall RMSE (see Figure 6). We then proceeded to standardize the model matrix, which had dimensions of 395 × 5. A preliminary Tikhonov regularization was then performed to obtain an initial guess for the mean coefficient vector ($\mu_{0}$) and its covariance matrix ($\Sigma_{0}$).The algorithm was then applied with the parameter update steps defined by Equations 4.4 and 4.5. Table 10 presents the estimated parameters from both the initial guess and the final iterative step of the warm-up estimation period.
@@
Date $SST$ $SSE$ $n$ $p$ $R^{2}$ $\hat{\sigma}$ $RMSE$
@@
The Individual values control chart in Figure 8 confirms that the residuals are randomly distributed with a zero mean. This interpretation is consistent with the residuals plot in 8(c). Additionally, the histogram in Figure 8 displays a symmetric distribution of residuals centered around zero.
@@
Date $SST$ $SSE$ $n$ $p$ $R^{2}$ $\hat{\sigma}$ $RMSE$
@@
Experiment 6
We applied an ensemble method, as discussed in section 5, to the previous experiment dataset. The same warm-up and real-time estimation procedures were followed. Nine distinct experts were employed, and the model provided by each expert is briefly described in Table 17.
The first expert used a model identical to that of the previous experiment, incorporating five independent variables (including an intercept), as detailed in Table 8. The remaining experts varied their models by including augmented dictionary terms. These augmentations involved either two-way interaction terms, squared terms, or the sine or cosine transformation of the existing variables.
Table 17 further illustrates the evolution of the weights distributed among the experts. At the initial iterative step, each expert was assigned an equal weight of $1/9=0.1111$. The updated weights at the end of the warm-up and real-time estimation stages demonstrate that the algorithm learned to adapt dynamically and prioritize expert contributions based on their performance. For example, at the final warm-up estimation step, there is no significant distinction in the weight distribution. However, by the final iterative steps of the real-time estimation, the weights of experts 2, 6 and 9 have been reduced to zero, likely because they committed larger prediction deviations compared to the other experts. Expert 9 received the highest weight allocation.
Table 17. Expert advice predictions models
Expert Interaction Squared Sine Cosine $p$ $t_{0}$ $t_{395}$ $t_{699}$ $t_{700}$
E1 No No No No 5 0.1111 0.1057 0.1732 0.1742
E2 Yes No No No 11 0.1111 0.1127 0.0000 0.0000
E3 No Yes No No 9 0.1111 0.1099 0.1472 0.1476
E4 No No Yes No 9 0.1111 0.1099 0.1988 0.1995
E5 No No No Yes 9 0.1111 0.1101 0.1757 0.1764
E6 Yes Yes No No 15 0.1111 0.1136 0.0000 0.0000
E7 Yes Yes Yes No 19 0.1111 0.1124 0.1005 0.0970
E8 Yes Yes Yes Yes 23 0.1111 0.1161 0.0000 0.0000
E9 No No Yes Yes 13 0.1111 0.1096 0.2047 0.2054
We do not present any additional figures or related performance metric tables here, as the overall results obtained from this ensemble method are almost identical to those of the previous experiment.
Experiment 7
In this experiment, we applied the randomized SINDy algorithm to the Elec2 dataset, a real-world binary classification problem. The objective of the dataset focuses on predicting the direction of electricity price changes (up or down) in New South Wales (NSW), Australia. Comprising 45,312 instances, the dataset includes five features and spans the period from May 7, 1996 to December 5, 1998. Table 18 provides a brief description of the variables and further details can be found in [Harries]. The dataset was sourced from Kaggle (https://www.kaggle.com/datasets/yashsharan/the-elec2-dataset). All features, with the exception of “day”, were pre-normalized to the range of 0 and 1 upon download. To maintain a uniform data scale, the “day” variable was subsequently normalized using min-max scaling. The class label distribution is shown in Table 19.
@@
Finally, the classification threshold was set to 0.5654, which was derived from the ROC curve (AUC 0.8313), in 12(a). The accuracy and other performance metrics are presented in 12(b). The ROC curve shows that our model has strong predictive power with high accuracy (above 80%).
@@
The Elec2 dataset is widely recognized as a benchmark for concept drift detection in the online learning literature. Following the Drift Detection Method (DDM) proposed by [Elec2Drift], the trace of model’s error rate, $p_{i}$, the probability of a false prediction up to step $i$), is presented in Figure 13. The left panel of Figure 13 shows no apparent major changes in the overall error rate. The right panel, a zoomed-in view focusing on the range of 0.12 to 0.16, reveals a slight increase in the error rate within an early window. However, after this period, the error rate consistently decreases and then stabilize, fluctuating around a level of 0.146 indicated by the red dashed line. This observed error trace differs from those presented in the study by [Elec2Drift] for the same dataset, which shows various moments of abrupt drifts. This discrepancy may be attributed to the dynamic and adaptive learning capabilities of our proposed model, highlighting its ability to adjust to changes in the data distribution even when resources for explicit major drift alerts are unavailable.
@@ 7. Discussion and Conclusion
In this paper, we propose a randomized sequential machine learning algorithm capable of handling dynamic data structures that does not rely on the classical iid assumption. Section 2 provides an introduction to the dynamic data arrival structure from which the algorithm was developed from a probabilistic perspective. Section 3 provides an extensive proof that mathematically demonstrates the proposed algorithm satisfies the PAC learning property. Additionally, a proof-of-concept was presented in the form of a theorem and a related proof. The proposed randomized algorithm considers a class of hypotheses or predictors that fulfill a probability density function. The learner takes action dynamically at every time step according to this density. More precisely, the learner uses the probability distribution of the predictors learned from the previous time step to make a prediction for the current time step and updates the density after the incoming data for the current time step is revealed. At each time step, the algorithm updates the weights assigned to the predictors using a gradient descent and a proximal algorithm, the latter being which is equivalent to a projection operator and ensures that the distributed weights form a valid probability density.
We were inspired by the SINDy (Sparse Identification of Nonlinear Dynamics) algorithm, a widely used data-driven modeling method for dynamics systems in recent years. We named our proposed algorithm “randomized SINDy”, which is presented in Section 4 and considers the possibility of augmenting independent variables. This augmentation is sometimes referred to as dictionary or feature library, resulting in a high-dimensional model matrix that admits a sparse representation. We introduced Tikhonov regularization to the objective function to overcome the problem of ill-conditioned data. Assuming the weights follow a multivariate normal distribution, the proximal step is unnecessary. Thus, the learner’s task is to estimate the parameters of the multivariate normal distribution. We explored two common applications, regression problem and binary classification with a logistic model, both of which include a regularized term in the objective function.
Section 5 details the application of the algorithm as an ensemble method, also known as the method of prediction with expert advice. We only considered the finite-dimensional case. Section 6 presents six experiments using simulated data and application on real-world data. The results of these experiments indicate the usefulness of our proposed algorithm.

### v2

@@ 2.2. Randomized SINDy for binary outcome
$\ell(y,z^{\top}\beta)=-\bigr[y\ln\left(h\left(z\right)\right)+\left(1-y\right)\ln\left(1-h\left(z\right)\right)\bigr],$
we note that there is no closed form for $\mathbb{E}\ell(y,z^{\top}\beta)$. Rather, we consider its approximation $\ell(y,z^{\top}\mu)$. Consequently, the objective function is as follows:
$\underbrace{\ell(y,z^{\top}\mu)+\lambda\left(\text{tr}\Sigma+\mu^{\top}\mu\right)}_{f_{1}}+\underbrace{1\!\!1_{p.s.d}(\Sigma)}_{f_{2}}$
@@
$\nabla_{\mu}f_{1}(\mu,\Sigma)=z\left(\sigma\left(z^{\top}\mu\right)-y\right)+2\lambda\mu,$
@@
$\mu_{t+1}=\mu_{t}-\eta_{t}\nabla_{\mu_{t}}f_{1}(\mu_{t},\Sigma_{t})=\mu_{t}-\eta_{t}\left(\mathbf{x}_{t}\left(\frac{1}{1+e^{-z^{\top}_{t}\mu_{t}}}-y_{t}\right)+2\lambda\mu_{t}\right),$
@@ 3. Experiments
In this section, we present some simulation results of models discussed in section 2. All experiments were executed using the R language.
@@
To evaluate the impact of the initial value, we vary $\mu_{0}$ between 0 and 1.8. As shown in Figure 2, the initial values influence the convergence rate. For a sequence observed at the 50th time step, the initial value of $\mu_{0}=0$ requires more iterations to achieve convergence.
@@
The first 250 time steps were designated as a warm-up period for model estimation. Within this warm-up period, we performed a rolling window cross-validation with 40 train-validation splits. The first training set consisted of 50 observations. For each subsequent split, the training set size was incremented by five observations (e.g.. 50, 55, …, up to 245 observations). Each corresponding validation set contained the five subsequent observations immediately following its training set. For each of these 40 train-validation splits, we fitted a regularized regression model. The regularization parameter ($\lambda$) was varied across 30 different values, ranging from 0 to 1. To determine the optimal $\lambda$, the minimum Root Mean Squared Error (RMSE) was calculated for each $\lambda$ value on each of the 40 validation sets. Subsequently, $\lambda=0.3448$ was selected as the optimal $\lambda$ as it yielded the lowest overall RMSE. Still within this warm-up period, we proceeded to standardize the model matrix, which has dimensions of $250\times 4$. A preliminary Tikhonov regularization was then performed to obtain an initial guess for the mean coefficient vector ($\mu_{0}$) and its covariance matrix ($\Sigma_{0})$. Subsequently, the algorithm was applied with the parameter update steps defined by Equations 2.5 and 2.6. The model parameters estimated at $t=250$ were used as the initial state for the subsequent time step. From $t=251$ onwards, each new incoming data dynamically updated both the standardization of the model matrix and the estimated parameters. With a learning rate of $\eta=1\times 10^{-3}$, Table 3 presents the estimated parameters at the initial and the final time step.
@@
$t$ SSTSST SSESSE $n$ $p$ $R^{2}$ $\hat{\sigma}$ RMSERMSE
@@
We simulated a binary sequence $Y\sim Ber(p)$, where $p$ was defined by the logistic model $P(Y=1\mid X)=\dfrac{1}{1+\exp\{-(1+2X_{1}+3X_{2})\}}.$ As in the previous experiment, a sequence of 500 time steps was generated. Initial logistic model parameters were obtained by fitting a ridge logistic regression with $\lambda=0$. The algorithm from Example 2 of section 2 was then applied with a fixed learning rate of $\eta=0.1$ . Table 5 shows the estimated coefficients at both the initial and final time steps.
@@
We applied the models discussed in section 2 to forecast the monthly U.S. unemployment rate. As described in the work of [8], the independent variables included are Initial Claims (IC) for unemployment insurance, the inflation rate (measured as the seasonal growth rate of the Consumer Price Index), and industrial growth (the seasonal growth rate of the Industrial Production Index). All data were sourced from the FRED database (https://fred.stlouisfed.org/). The analysis period spans from January 1967 to May 2025. Table 7 gives a brief description of the variables used for forecasting. While most variables are originally monthly, the weekly Initial Claims data was averaged to a monthly frequency.
@@
We used the initial 30 years of data, spanning from January 1967 to December 1999, as a warm-up period for our online learning model. Based on a rolling expanding window approach, we began with 20 years of data for training to make a one-year-ahead prediction. This process generated 12 distinct train-validation set splits for model evaluation. For each of these 12 training sets, we performed regression with Tikhonov regularization, exploring 30 different regularization parameter ($\lambda$) values ranging from 0 to 1. The optimal regularization term was identified as being the value of 0, as this produced the minimum overall RMSE (see Figure 4). We then proceeded to standardize the model matrix, which had dimensions of 395 × 5. A preliminary Tikhonov regularization was then performed to obtain an initial guess for the mean coefficient vector ($\mu_{0}$) and its covariance matrix ($\Sigma_{0}$).The algorithm was then applied with the parameter update steps defined by Equations 2.5 and 2.6. Table 9 presents the estimated parameters from both the initial guess and the final iterative step of the warm-up estimation period.
@@
Date SSTSST SSESSE $n$ $p$ $R^{2}$ $\hat{\sigma}$ RMSERMSE
@@
The Individual values control chart in Figure 6 confirms that the residuals are randomly distributed with a zero mean. This interpretation is consistent with the residuals plot in 6(c). Additionally, the histogram in Figure 6 displays a symmetric distribution of residuals centered around zero.
@@
Date SSTSST SSESSE $n$ $p$ $R^{2}$ $\hat{\sigma}$ RMSERMSE
@@
Experiment 5
In this experiment, we applied the randomized SINDy algorithm to the Elec2 dataset, a real-world binary classification problem. The objective of the dataset focuses on predicting the direction of electricity price changes (up or down) in New South Wales (NSW), Australia. Comprising 45,312 instances, the dataset includes five features and spans the period from May 7, 1996 to December 5, 1998. Table 16 provides a brief description of the variables and further details can be found in [7]. The dataset was sourced from Kaggle (https://www.kaggle.com/datasets/yashsharan/the-elec2-dataset). All features, with the exception of “day”, were pre-normalized to the range of 0 and 1 upon download. To maintain a uniform data scale, the “day” variable was subsequently normalized using min-max scaling. The class label distribution is shown in Table 17.
@@
Finally, the classification threshold was set to 0.5654, which was derived from the ROC curve (AUC 0.8313), in 10(a). The accuracy and other performance metrics are presented in 10(b). The ROC curve shows that our model has strong predictive power with high accuracy (above 80%).
@@
The Elec2 dataset is widely recognized as a benchmark for concept drift detection in the online learning literature. Following the Drift Detection Method (DDM) proposed by [6], the trace of model’s error rate, $p_{i}$, the probability of a false prediction up to step $i$), is presented in Figure 11. The left panel of Figure 11 shows no apparent major changes in the overall error rate. The right panel, a zoomed-in view focusing on the range of 0.12 to 0.16, reveals a slight increase in the error rate within an early window. However, after this period, the error rate consistently decreases and then stabilize, fluctuating around a level of 0.146 indicated by the red dashed line. This observed error trace differs from those presented in the study by [6] for the same dataset, which shows various moments of abrupt drifts. This discrepancy may be attributed to the dynamic and adaptive learning capabilities of our proposed model, highlighting its ability to adjust to changes in the data distribution even when resources for explicit major drift alerts are unavailable.
@@ 4. Discussion and Conclusion
In this paper, we propose a randomized sequential machine learning algorithm capable of handling dynamic data structures that does not rely on the classical iid assumption. Section 2 provides an introduction to the dynamic data arrival structure from which the algorithm was developed from a probabilistic perspective. We also provide in this section an extensive proof that mathematically demonstrates the proposed algorithm satisfies the PAC learning property. Additionally, a proof-of-concept was presented in the form of a theorem and a related proof. The proposed randomized algorithm considers a class of hypotheses or predictors that fulfill a probability density function. The learner takes action dynamically at every time step according to this density. More precisely, the learner uses the probability distribution of the predictors learned from the previous time step to make a prediction for the current time step and updates the density after the incoming data for the current time step is revealed. At each time step, the algorithm updates the weights assigned to the predictors using a gradient descent and a proximal algorithm, the latter being which is equivalent to a projection operator and ensures that the distributed weights form a valid probability density.
We were inspired by the SINDy (Sparse Identification of Nonlinear Dynamics) algorithm, a widely used data-driven modeling method for dynamics systems in recent years. We named our proposed algorithm “randomized SINDy” which considers the possibility of augmenting independent variables. This augmentation is sometimes referred to as dictionary or feature library, resulting in a high-dimensional model matrix that admits a sparse representation. We introduced Tikhonov regularization to the objective function to overcome the problem of ill-conditioned data. Assuming the weights follow a multivariate normal distribution, the proximal step is unnecessary. Thus, the learner’s task is to estimate the parameters of the multivariate normal distribution. We explored two common applications, regression problem and binary classification with a logistic model, both of which include a regularized term in the objective function.
Section 3 presents five experiments using simulated data and application on real-world data. The results of these experiments indicate the usefulness of our proposed algorithm.